OpenStudy (anonymous):

Determine the area of a circle that has a radius of 8 cm. If necessary round answers to nearest tenth.

OpenStudy (anonymous):

Hey! So, yeah, there's this really cool fact about circles that lets you find the area just by knowing the radius! All you have to do is \[\Large \pi*r^2\]

OpenStudy (da_scienceman):

tbrroks pls substitute as Smoothmath said

OpenStudy (anonymous):

Here's a quick example. If I have this circle:  I notice that the radius is 2, so to find the area, all I have to do is: \[\Large \pi*(2)^2 = \pi*4 = 4\pi \approx 12.566\]

OpenStudy (anonymous):

thanks for given me a simplified example it helps alot for my other questions .

OpenStudy (anonymous):

but if it say round to nearest tenth would i make it 12.57 ?

OpenStudy (anonymous):